angelica and her brother, morton, had a total fo $100 they earned washing walls. they heard about a fantastic garage sale and went to see what they could find. when they saw the prices for paperback books, jeans, and winter jackets, they decided to buy 100 items for their $100. if hte jackets were $9 each, the jeans $3 a pair, and the books 5 for $1, how many of each did they buy?

Answer by stanbon(75887)

they decided to buy 100 items for their $100. if hte jackets were $9 each, the jeans $3 a pair, and the books 5 for $1, how many of each did they buy?
----------------
Let # of jackets be "k" ; Value of those jackets is 900k cents
Let # of jeans be "n" ; value of those jeans is 300n cents
Let # of books be "b" ; value of those books is 20n cents
-------------
EQUATIONS:
k + n + b = 100
900k + 300n + 20b = 10000 cents
--------
k + n + b = 100
45k + 15n +b = 500
---------
Subtract 1st from 2nd to get:
44k + 13n = 400
---------------
Make a list of the multiples of 44 and another list of the multiples
of 13. Try to find a sum that equals 400. That should tell you
what k and n are. Using that information you can find b.
It can be done because k and n must be Natural or Counting numbers.
--------------
Cheers,
Stan H.